OpenStudy (anonymous):

intergrate (x)cube/(x)sq - 1

OpenStudy (anonymous):

\[\int\limits_{}^{} x^{3} / (x^{2} - 1)\]

OpenStudy (lgbasallote):

\[\int \frac{x^3}{x^2 - 1}\] divide this thing first x - x/x^2 - 1 ________ x^2 - 1 | x^3 x^3 ------- 0 - x now sub \[\int (x - \frac{x}{x^2 - 1} )dx\] distribute the integral \[\int xdx - \int \frac{x}{x^2 - 1}dx\] \[\int xdx - \frac 12 \int \frac{1}{u - 1}du\]

OpenStudy (anonymous):

can anyone explain the 2nd part to the solution provided by lgbastalloe?

OpenStudy (lgbasallote):

division of polynomials?

OpenStudy (lgbasallote):

or the substitution?

OpenStudy (anonymous):

the substitution

OpenStudy (lgbasallote):

you got the division though right?

OpenStudy (anonymous):

yep i got it. only the substituition.

OpenStudy (lgbasallote):

if i had \[\int \frac{4}{2}dx\]then i can just write this as \[\int 2dx\] right? since the quotient of 4 divided by 2 is 2 same here the quotient of \[\frac{x^3}{x^2 - 1} \implies x - \frac{x}{x^2 - 1}\] make sense?

OpenStudy (anonymous):

yes. i get that. but how did you substitute U into the equation?

OpenStudy (lgbasallote):

ohh u...

OpenStudy (anonymous):

when i did it, i got stuck. i couldnt get the same outcome as you.

OpenStudy (lgbasallote):

\[\int \frac{x}{x^2 - 1}dx\] let u = x^2 - 1 du = 2x dx du/2 = xdx \[\int \frac{x}{x^2 - 1}dx \implies \frac{1}{2} \int \frac{du}{u}\] lol just noticed i made a mistake in what i write a while ago

OpenStudy (anonymous):

OH Thanks alot. i get it now.

OpenStudy (lgbasallote):

\[\huge \color{darkblue}{\textbf{<tips hat>}}\]